# Vector-Quantized Vision Foundation Models for Object-Centric Learning

### v3

@@ Abstract.
Perceiving visual scenes as objects and background—like humans do—Object-Centric Learning (OCL) aggregates image or video feature maps into object-level feature vectors, termed slots. OCL’s self-supervision of reconstructing the input from these aggregated slots struggles with complex object textures, thus Vision Foundation Model (VFM) representations are used as the aggregation input and reconstruction target. However, existing methods leverage VFM representations in diverse ways and often fail to fully exploit their potential. In response, we propose a clean architecture—Vector-Quantized VFMs for OCL (VQ-VFM-OCL, or VVO)—that unifies mainstream OCL methods. The key to our unification is simple yet effective, just shared quantizing the same VFM representation as the reconstruction target. Through mathematical modeling and statistical verification, we further analyze why VFM representations facilitate OCL aggregation and how their shared quantization as reconstruction targets strengthens OCL supervision. Experiments show that across different VFMs, aggregators and decoders, our VVO consistently outperforms baselines in object discovery and recognition, as well as downstream visual prediction and reasoning. The implementation and model checkpoints are available on https://github.com/Genera1Z/VQ-VFM-OCL.
isbn: 978-1-4503-XXXX-X/2018/06submissionid: 7307copyright: none
@@ 1. Introduction
Objects can form highly diverse visual scenes through arrangements and combinations. But mainstream methods based on feature patches or a single feature vector disregard such compositionality. Inspired by human vision cognition (Bar, 2004; Cavanagh, 2011; Palmeri and Gauthier, 2004), Object-Centric Learning (OCL) decomposes visual scenes into multiple feature vectors, known as slots, each corresponding to an object or the background, thus enabling improved modeling of relationships and dynamics among objects. Object-centric representations have demonstrated superiority in advanced vision tasks, like prediction, reasoning, planning, and decision-making (Wu et al., 2023a), as well as in interactions between visual modality and other modalities (Zhang et al., 2024; Wang et al., 2024b).
@@
OCL relies on pixel textures to discover objects. The early milestone (Locatello et al., 2020) reconstructs input pixels for supervision, which usually fails on realistic objects. Some (Kipf et al., 2022; Elsayed et al., 2022) reconstruct optical flow or depth map to mitigate textural noises, but at the cost of expensive annotations. Some (Singh et al., 2022a, c) reconstruct input’s VAE (Variational Autoencoder) representation, where super-pixels are codebook codes shared across samples, thus suppressing pixel redundancy and facilitating aggregation from the feature map into slots. Recent advances (Seitzer et al., 2023; Wu et al., 2023b) leverage Vision Foundation Models (VFMs) (Caron et al., 2021; Oquab et al., 2023) to extract feature maps with better object separability, significantly improving OCL performance.
@@
To address these issues, we propose a clean architecture—Vector-Quantized VFMs for OCL (VQ-VFM-OCL, or VVO)—that unifies mainstream OCL methods. In this architecture, as shown in Figure 2, our VVO supports different VFMs for encoding, different OCL aggregators and different OCL decoders. The key to such unification is very simple—We quantize the representations from the same VFM, rather than from another encoder, as the reconstruction target. We also mathematically and statistically analyze why VFM representations facilitate OCL aggregation and how their shared quantization as reconstruction targets strengthens OCL supervision.
@@ 2. Related Work
SSseg vs OCL. Both Self-Supervised Segmentation (SSseg) (Ge et al., 2023; Wang et al., 2024a) and OCL can discovery objects, i.e., segmenting objects and the background in images or video frames without supervision. SSseg only outputs object segmentation masks, while OCL further represents objects as slots. These slots can be directly used in downstream object-centric world models (Wu et al., 2023a) for dynamics modeling, which is demanded in advanced vision tasks. Thus a direct comparison between them is beyond the scope of this work.
OCL encoding. The stronger encoder that extracts feature maps of better objectness contributes more to OCL performance. Early milestone methods like IODINE(Greff et al., 2019) and SA (Locatello et al., 2020) use small naive CNNs (Krizhevsky et al., 2012) as OCL encoder. Followups like SAVi (Kipf et al., 2022), SAVi++ (Elsayed et al., 2022), SLATE (Singh et al., 2022a) and STEVE (Singh et al., 2022c) employ pretrained ResNets (He et al., 2016), and fine-tune them on OCL datasets. State-of-the-arts like SlotDiffusion (Wu et al., 2023b) and DINOSAUR (Seitzer et al., 2023) utilize VFMs like DINO (Caron et al., 2021) and DINO2 (Oquab et al., 2023) ViTs (Vision Transformers) to extract highly object-separable feature map from input pixels, improving OCL performance significantly. Although SAM (Kirillov et al., 2023) and SAM2 (Ravi et al., 2024) are also well recognized VFMs, they remain unexploited in OCL setting. Our VVO supports various VFMs for OCL encoding.
@@ 3.1. Unify OCL
As shown in Figure 2, our method adopts an architectural design of encoder-aggregator/quantizer-decoder.
Firstly, OCL encoder $\bm{\phi}_{\mathrm{e}}$ transforms an input image or video frame $\bm{X}\in\mathbb{R}^{h_{0}\times w_{0}\times c_{0}}$, which is an observation of some visual scene, into a dense feature map $\bm{Z}\in\mathbb{R}^{h\times w\times c}$, for the following query-based aggregation:
@@
where $\bm{\phi}_{\mathrm{e}}$ can be parameterized as pretrained frozen VFMs, like DINO (Caron et al., 2021), DINO2 (Oquab et al., 2023), SAM (Kirillov et al., 2023) and SAM2 (Ravi et al., 2024). As OCL relies on textures to separate objects, $\bm{\phi}_{\mathrm{e}}$ should handle complex textures of objects, making VFMs necessary here. We will detail the reasons in Section 3.2.
Secondly, given queries $\bm{S}_{0}\in\mathbb{R}^{n\times c}$, OCL aggregator $\bm{\phi}_{\mathrm{a}}$ transforms $\bm{Z}$ into multiple feature vectors or slots $\bm{S}\in\mathbb{R}^{n\times c}$ and byproduct segmentation masks $\bm{M}\in\mathbb{R}^{h\times w}$, corresponding to objects or the background in the scene:
@@
For video OCL, there is a recurrent module turning current $\bm{S}$ into new queries $\bm{S}_{0}$ for the next time step. Such module can be Transformer encoder blocks (Vaswani et al., 2017).
@@
where $\bm{\phi}_{\mathrm{q}}$ can be parameterized as some Vector Quantization (VQ) module (Im Im et al., 2017; Van Den Oord et al., 2017). But we meticulously design our own VQ variant, as detailed in Section 3.2. $\bm{\phi}_{\mathrm{q}}$ is pretrained in VAE framework and is frozen afterwards, where the encoder is shared from the frozen OCL encoder $\bm{\phi}_{\mathrm{e}}$, and the decoder is a typical VAE decoder. We will explain why not use a separate typical VAE encoder in Section 3.2.
Thirdly, OCL decoder $\bm{\phi}_{\mathrm{d}}$ transforms $\bm{S}$ into reconstruction $\bm{Q}^{\prime}\in\mathbb{R}^{h\times w\times c}$ with destructed $\bm{Q}$ as the condition:
@@ 3.2. Utilize VFMs in OCL
As shown in Figure 1, VFM representations have better objectness than non-VFMs, even under naive kMeans clustering [^1] https://scikit-learn.org/stable/modules/generated/sklearn.cluster.KMeans.html. OCL aggregation is essentially a clustering whose initial centroids are trainable (Jia et al., 2023). Thus, we expect $\bm{\phi}_{\mathrm{a}}$ to aggregate VFM’s $\bm{Z}$ into slots $\bm{S}$ better under queries $\bm{S}_{0}$. Previous methods like DINOSAUR (Seitzer et al., 2023) have already exploited this but without any reason. In Section 5.1, we analyze this both theoretically and empirically.
@@
Given the direct VFM feature map $\bm{Z}$, we transform it with a naive CNN to eliminate the positional information for better quantization. We quantize the adjusted $\bm{Z}$ using our VQ variant as $\bm{\phi}_{\mathrm{q}}$ to obtain the reconstruction target $\bm{Q}$.
@@
(7) $\bm{D}=||\bm{Z}-\bm{T}||_{2}^{2}$
where $\mathrm{sg}(\cdot)$ is stop-gradient; $\bm{T}\in\mathbb{R}^{m\times c}$ is the codebook to be used for quantizing $\bm{Z}$; $\bm{D}\in\mathbb{R}^{h\times w\times m}$ is the matching distance between every super-pixel and every code.
We convert distances to probabilities and select the most matched codes to form the quantization $\bm{Q}$:
@@
where $\bm{G}\in\mathbb{R}^{h\times w\times m}$ is the Gumbel noise and $\tau$ is the temperature. We devise annealing residual connection (Zhao et al., 2024a, b) after Equation 10 to stabilize pre-training:
@@
where $\alpha$ is scheduled from 1 to 0 during pre-training using cosine-annealing. Besides typical losses of reconstruction, alignment and commitment (Van Den Oord et al., 2017), we regularize the adjusted $\bm{Z}$’s elements to be normalized:
(13) $l_{n}=\lambda\mathrm{MSE}(\bm{Z},\mathrm{sg}(\frac{\bm{Z}-\mathbb{E}[\bm{Z}]}{\sqrt{\mathbb{V}[\bm{Z}]+\epsilon}}))$
@@
With all samples’ feature maps $\bm{Z}$ being represented with one codebook $\bm{T}$, the quantization $\bm{Q}$ naturally gains cross-sample consistency and suppresses information redundancy or noise. This helps the aggregation with queries $\bm{S}_{0}$, which are also shared across samples. After such tokenization, our reconstruction target can be either regressed using mixture and diffusion decoding or classified using auto-regressive decoding. In contrast, methods like SLATE (Singh et al., 2022a) and SlotDiffusion (Wu et al., 2023b) use separate VAE and OCL encoders, which as shown in Figure 1 causes a distribution gap between $\bm{Q}$ and $\bm{Z}$. Thus, we expect shared VFM representation quantization as reconstruction targets to strengthen OCL supervision. In Section 5.2, we analyze this both theoretically and empirically.
3.3. Compare Architectures
@@
As shown in Figure 3, we compare the baselines with our VVO in a unified perspective. Specifically,
- Our VVO: (1) VFMs are employed for OCL encoding and their features are fed to the aggregator directly, which eases OCL aggregation, as formulated in Section 3.2. (2) VFM features are shared quantized as reconstruction targets, which strengthens OCL supervision, as formulated in Section 3.2.
- SLATE0 (the official version) (Singh et al., 2022a): (1) VFM features are NOT directly fed to the aggregator. (2) VFM features are shared discretized to scalar numbers and re-embedded into features to be learned latter. This loses much information of VFM features, while shared quantization can preserve such information (Van Den Oord et al., 2017).
- SLATE (the unofficially improved version; we adopt it) (Jia et al., 2023; Singh et al., 2022c; Wu et al., 2023b) / STEVE (Singh et al., 2022c): (1) Same as our VVO. (2) Reconstruction targets are discretized from features of a separate VAE encoder, instead of quantized from features shared from the same VFM for OCL encoding, which causes optimization noises.
- SlotDiffusion (Wu et al., 2023b) / LSD (Jiang et al., 2023): (1) Same as our VVO. (2) Reconstruction targets are quantized from features of a separate VAE encoder, instead of sharing the same VFM for OCL encoding, which causes optimization noises.
- DINOSAUR (Seitzer et al., 2023) / VideoSAUR (Zadaianchuk et al., 2024) / SPOT (Kakogeorgiou et al., 2024): (1) Same as our VVO. (2) Reconstruction targets are shared from the same VFM for OCL encoding but without quantization, which causes optimization noises.
Our VVO adopts a clean and unified architecture with the above-mentioned two key designs. The first design, direct VFM feature extraction for OCL aggregation, is not proposed by us but we are the first to provide both theoretical and empirical analyses. The second design, shared VFM feature quantization as reconstruction targets, is proposed by us. Such shared quantization is also the key point that our VVO supports various OCL decoders. Specifically,
- Mixture-based OCL decoders, e.g., CNN (Locatello et al., 2020; Kipf et al., 2022; Elsayed et al., 2022), MLP (Seitzer et al., 2023) and SlotMixer (Zadaianchuk et al., 2024), are originally designed for continuous features. As quantized features are actually the clustering centroids of continuous features, these decoders are inherently compatible with our shared quantized VFM features.
- Auto-regressive OCL decoders, e.g., the Transformer decoder (Singh et al., 2022a, c) and Transformer9 (Kakogeorgiou et al., 2024), are originally designed for discretized features, while (Seitzer et al., 2023) and (Kakogeorgiou et al., 2024) also explore their applicability to continuous features. Thus these decoders are naturally applicable to our shared quantized VFM features.
- Diffusion-based OCl decoders, e.g., Conditional Diffusion (Wu et al., 2023b; Jiang et al., 2023), are originally designed for dicretized features. Thus if without our shared quantized VFM features as reconstruction targets, then this type of OCL decoders would not be unified into our architecture.
@@
256$\times$256 (224) ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU
@@
256$\times$256 ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU

### v5

@@ Abstract.
Object-Centric Learning (OCL) aggregates image or video feature maps into object-level feature vectors, termed slots. It’s self-supervision of reconstructing the input from slots struggles with complex object textures, thus Vision Foundation Model (VFM) representations are used as the aggregation input and reconstruction target. Existing methods leverage VFM representations in diverse ways yet fail to fully exploit their potential. In response, we propose a unified architecture, Vector-Quantized VFMs for OCL (VQ-VFM-OCL, or VVO). The key to our unification is simply shared quantizing VFM representations in OCL aggregation and decoding. Experiments show that across different VFMs, aggregators and decoders, our VVO consistently outperforms baselines in object discovery and recognition, as well as downstream visual prediction and reasoning. We also mathematically analyze why VFM representations facilitate OCL aggregation and why their shared quantization as reconstruction targets strengthens OCL supervision. Our source code and model checkpoints are available on https://github.com/Genera1Z/VQ-VFM-OCL.
Object-Centric Learning, Vision Foundation Model, Vector Quantization, Object Representation, Visual Prediction, Visual Reasoning
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3754566isbn: 979-8-4007-2035-2/2025/10submissionid: mfp7307copyright: noneccs: Computing methodologies Computer vision representations
@@ 1. Introduction
Objects can form highly diverse visual scenes through arrangements and combinations. But mainstream methods based on feature patches or a single feature vector disregard such compositionality. Inspired by human vision cognition (Bar, 2004; Cavanagh, 2011; Palmeri and Gauthier, 2004), Object-Centric Learning (OCL) decomposes visual scenes into multiple feature vectors, known as slots, each corresponding to an object or the background, thus enabling improved modeling of relationships and dynamics among objects. Object-centric representations have demonstrated superiority in advanced vision tasks, like prediction, reasoning, planning, and decision-making (Wu et al., 2023a), as well as in interactions between visual modality and other modalities (Zhang et al., 2024; Wang et al., 2024).
@@
OCL relies on pixel textures to discover objects. The early milestone (Locatello et al., 2020) reconstructs input pixels as supervision, usually failing on realistic objects. Some (Kipf et al., 2022; Elsayed et al., 2022) reconstruct optical flow or depth map to mitigate textural noises, at the cost of expensive annotations. Some (Singh et al., 2022a, c) reconstruct input’s VAE (Variational Autoencoder) representation, whose super-pixels are codebook codes, thus suppressing pixel redundancy and facilitating aggregation from features into slots. Recent advances (Seitzer et al., 2023; Wu et al., 2023b) use Vision Foundation Models (VFMs) (Caron et al., 2021; Oquab et al., 2023) to extract features with better object separability, boosting OCL significantly.
@@
To address these issues, we propose a clean architecture–Vector-Quantized VFMs for OCL (VQ-VFM-OCL, or VVO)–that unifies mainstream OCL methods. In this architecture, as shown in Figure 2, our VVO supports different VFMs for encoding, different OCL aggregators and different OCL decoders. The key to such unification is very simple–We quantize the representations from the same VFM, rather than from another encoder, as the reconstruction target. We also mathematically analyze why VFM representations facilitate OCL aggregation and why their shared quantization as reconstruction targets strengthens OCL supervision.
@@ 2. Related Work
OCL encoding. Early milestone methods like IODINE(Greff et al., 2019) and SA (Locatello et al., 2020) use small naive CNNs (Krizhevsky et al., 2012) as OCL encoder. Followups like SAVi (Kipf et al., 2022), SAVi++ (Elsayed et al., 2022), SLATE (Singh et al., 2022a) and STEVE (Singh et al., 2022c) employ pretrained ResNets (He et al., 2016), and fine-tune them on OCL datasets. State-of-the-arts like SlotDiffusion (Wu et al., 2023b) and DINOSAUR (Seitzer et al., 2023) utilize VFMs like DINO (Caron et al., 2021) and DINO2 (Oquab et al., 2023) ViTs (Vision Transformers) to extract highly object-separable feature map from input pixels, improving OCL performance significantly. SAM (Kirillov et al., 2023) and SAM2 (Ravi et al., 2024) are also well recognized VFMs yet remain unexploited in OCL setting. Our VVO supports various VFMs for OCL encoding.
@@ 3.1. Unify OCL
Our method adopts an architectural design as shown in Figure 2.
Firstly, OCL encoder $\bm{\phi}_{\mathrm{e}}$ transforms input, an image or video frame $\bm{X}\in\mathbb{R}^{h_{0}\times w_{0}\times c_{0}}$ of some visual scene, into a dense feature map $\bm{Z}\in\mathbb{R}^{h\times w\times c}$, for the following query-based aggregation:
@@
where $\bm{\phi}_{\mathrm{e}}$ can be parameterized as pretrained VFMs, like DINO (Caron et al., 2021), DINO2 (Oquab et al., 2023), SAM (Kirillov et al., 2023) and SAM2 (Ravi et al., 2024). As OCL relies on textures to separate objects, $\bm{\phi}_{\mathrm{e}}$ should handle complex textures of objects, making VFMs necessary here. We will explain it in Section 3.2.
Secondly, given queries $\bm{S}_{0}\in\mathbb{R}^{n\times c}$, OCL aggregator $\bm{\phi}_{\mathrm{a}}$ transforms $\bm{Z}$ into multiple feature vectors or slots $\bm{S}\in\mathbb{R}^{n\times c}$ and corresponding byproduct segmentation masks $\bm{M}\in\mathbb{R}^{h\times w}$, each representing a specific object or background in the scene:
@@
For video OCL, there is a recurrent module transitioning current slots $\bm{S}$ into new queries $\bm{S}_{0}$ for the next time step. Such module can be parameterized as a Transformer encoder block (Vaswani et al., 2017).
@@
where $\bm{\phi}_{\mathrm{q}}$ can be parameterized as some Vector Quantization (VQ) (Im Im et al., 2017; Van Den Oord et al., 2017). But we meticulously design our own VQ variant, as detailed in Section 3.2. $\bm{\phi}_{\mathrm{q}}$ is pretrained in VAE framework and is frozen afterwards, where the encoder is shared from the frozen OCL encoder $\bm{\phi}_{\mathrm{e}}$, and the decoder is a typical VAE decoder. We will explain why not use a separate typical VAE encoder in Section 3.2.
Thirdly, OCL decoder $\bm{\phi}_{\mathrm{d}}$ transforms slots $\bm{S}$ into reconstruction $\bm{Q}^{\prime}\in\mathbb{R}^{h\times w\times c}$ with destructed $\bm{Q}$ as the condition:
@@ 3.2. Utilize VFMs in OCL
As shown in Figure 1, VFM representations have better objectness than non-VFMs, even under naive kMeans clustering [^1] https://scikit-learn.org/stable/modules/generated/sklearn.cluster.KMeans.html. OCL aggregation is essentially a clustering whose initial centroids are trainable (Jia et al., 2023). Thus, we expect $\bm{\phi}_{\mathrm{a}}$ to aggregate VFM’s $\bm{Z}$ into slots $\bm{S}$ better under queries $\bm{S}_{0}$. Previous methods like DINOSAUR (Seitzer et al., 2023) have already exploited this but without any reason.
@@
Given the VFM feature $\bm{Z}$, we adjust it via a small CNN to eliminate the positional information for better quantization; Then quantize it using our VQ variant $\bm{\phi}_{\mathrm{q}}$ as the reconstruction target $\bm{Q}$.
@@
[…]
where $\mathrm{sg}(\cdot)$ is stop-gradient; $\bm{T}\in\mathbb{R}^{m\times c}$ is the codebook for quantizing $\bm{Z}$; $\bm{D}\in\mathbb{R}^{h\times w\times m}$ is the matching distance between every super-pixel in $\bm{Z}$ and every code in $\bm{T}$.
We convert distances to probabilities and select the most matched codes to form the quantization $\bm{Q}$ as the reconstruction target:
@@
where $\bm{G}\in\mathbb{R}^{h\times w\times m}$ is Gumbel noise and $\tau$ is the temperature. Training-time annealing residual connection (Zhao et al., 2024a, b) is added after Equation 10 to stabilize pre-training:
@@
where $\alpha$ is scheduled from 1 to 0 during pre-training using cosine-annealing. Besides typical losses of reconstruction, alignment and commitment (Van Den Oord et al., 2017), we regularize the adjusted $\bm{Z}$ to be normal:
(13) $l_{\mathrm{n}}=\lambda\mathrm{MSE}(\bm{Z},\mathrm{sg}(\frac{\bm{Z}-\mathbb{E}[\bm{Z}]}{\sqrt{\mathbb{V}[\bm{Z}]+\epsilon}}))$
@@
With all samples’ feature maps $\bm{Z}$ being represented with one codebook $\bm{T}$, the quantization $\bm{Q}$ naturally gains cross-sample consistency, helping the aggregation with queries $\bm{S}_{0}$, which are also shared across samples. Such tokenization is compatible with both regression and classification decoding. In contrast, methods like SLATE (Singh et al., 2022a) and SlotDiffusion (Wu et al., 2023b) are faced with distribution gaps between $\bm{Q}$ and $\bm{Z}$, shown in Figure 1, due to separate VAE and OCL encoders. Thus, we expect shared VFM representation quantization as reconstruction targets to strengthen OCL supervision.
@@
$\times256$ (224) ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU
@@
$\times256$ ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU
@@
3.3. Compare Architectures
As shown in Figure 3, we compare baselines methods with our VVO in a unified perspective. Specifically,
- Our VVO: (1) VFMs are employed for OCL encoding and their features are fed to the aggregator directly, which eases OCL aggregation, as formulated in Section 3.2. (2) VFM features are shared quantized as reconstruction targets, which strengthens OCL self supervision, as formulated in Section 3.2.
- SLATE0 (the official version) (Singh et al., 2022a): (1) VFM features are NOT directly fed to the aggregator. (2) VFM features are shared discretized to scalar numbers and re-embedded into features to be learned latter. This loses much information of VFM features.
- SLATE (the improved version; adopted here) / STEVE (Jia et al., 2023; Singh et al., 2022c; Wu et al., 2023b): (1) Same as VVO. (2) Reconstruction targets are discretized from features of separate VAE encoding, not quantized from features shared from OCL encoding, causing optimization noises.
- SlotDiffusion (Wu et al., 2023b) / LSD (Jiang et al., 2023): (1) Same as VVO. (2) Reconstruction targets are quantized from separate VAE encoding features, not sharing OCL encoding features, causing optimization noises.
- DINOSAUR (Seitzer et al., 2023) / VideoSAUR (Zadaianchuk et al., 2024) / SPOT (Kakogeorgiou et al., 2024): (1) Same as VVO. (2) Reconstruction targets are shared from VFM features of OCL encoding without quantization, causing optimization noises.
Our VVO realizes a clean and unified architecture based on the above-mentioned two key designs. Specifically,
- Mixture-based OCL decoders, e.g., CNN (Locatello et al., 2020; Kipf et al., 2022; Elsayed et al., 2022), MLP (Seitzer et al., 2023) and SlotMixer (Zadaianchuk et al., 2024), are originally designed for continuous features, thus are compatible with our shared quantized VFM features.
- Auto-regressive OCL decoders, e.g., the Transformer decoder (Singh et al., 2022a, c) and Transformer9 (Kakogeorgiou et al., 2024), are designed for discretized features, while also showing applicability to continuous features (Seitzer et al., 2023; Kakogeorgiou et al., 2024). Thus they are applicable to our shared quantized VFM features.
- Diffusion-based OCl decoders, e.g., conditional Diffusion (Wu et al., 2023b; Jiang et al., 2023), work on low-dimension features, necessitating our shared quantization on the continuous high-dimensional VFMs features.